OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)?

OpenStudy (anonymous):

use two point formula in finding the equation of the line. .

hero (hero):

\[(x_1, y_1) = (-2,1)\]\[(x_2, y_2) = (4,13)\]\[m = \frac{y_2 - y_1}{x_2 - x_1}\] Substitute those points into the slope formula to find m, the slope.

hero (hero):

Once you have found m, then use this point-slope formula: \[(y-y_1) = m(x-x_1)\] Once you have inserted the points, along with the value for m, you're done.

OpenStudy (anonymous):

@hero given the m (slope) formula, substitute it to the the point sole form to find the equation og the line given two points here...  use the formula. .